wiki:UC20190603

Version 1 (modified by trac, 6 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 6/03/19

Attending: Nancy Chanover (NMSU), Mark Klaene (APO), Bill Ketzeback (APO), Kevin Schlaufman (JHU), Misty Bentz (GSU), Joanne Hughes (Seattle U), Ben Williams (UW), Mike Skrutskie (UVa), Sarah Tuttle (UW), John Wisniewski (OU), Russet McMillan (APO), John Bally (CU)


Nancy started by welcoming a new Users Committee member - Kevin Schlaufman is replacing MaryBeth Kaiser as the institutional representative from Johns Hopkins University. Welcome, Kevin!

User feedback and comments from institutional representatives:

  • Seattle - no report
  • NAPG - no report
  • FGCU - no report
  • Washington - no report
  • Wyoming - no report
  • UVa - nothing to report
  • Oklahoma - nothing to report
  • NMSU - Moire reported a filter placement issue with a new filter that she purchased for use with ARCTIC and wanted to know whether others were experiencing similar problems. Russet said that she spent some engineering time on this issue. There have been a couple of confirmed problems and a couple of times where confusion was the issue (and not a mechanical problem). One of the confirmed problems was with Moire’s filter in a particular wheel; when they tested the wheel and the filter (in a different wheel) separately they both worked fine. So although they haven’t been able to replicate the problem, they do have a workaround for the time being (i.e., use Moire's filter in the wheel in which it works!).
  • JHU - nothing to report
  • Georgia State - nothing to report
  • Colorado - nothing to report

Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 5/02/2019 - 5/29/2019

0) Overview

The site has experienced some atypical spring weather with cooler temperatures and some moisture, combined with more typical spring winds. Forest moisture is good and currently there is only a moderate fire danger. Observing has been fairly productive with weather and systems cooperating. This month was a busy time on the mountain with a UVa class trip and observing runs with NAIC and PAO, two visiting instruments. Deanna Naugle, who handled most of the housing reservations, has resigned and Tina Blay (new hire) and Frances Cope are handling most of these functions now. We also held a nice luncheon to say goodbye to NMSU faculty member and former ARC Board Chair Rene Walterbos, who is retiring, along with Deanna and Jon Brinkmann, who retired in January.

1) Telescope

The telescope has been performing nominally.

2) Instruments

DIS: The DIS blue camera continues to see contamination issues reappear rapidly after servicing, while the red camera contamination appears more slowly.

TripleSpec is operational.

Agile is operational

Echelle is operational. The new Ecam is in place and operable. Documentation is being updated.

NIC-FPS is operational.

ARCTIC is operational; we are working on some minor filter issues.

3) ARCSAT

Operating nominally. We await direction on the future of ARCSAT from the Strategic Planning meeting in June.


Additional telescope and instrument discussion:

The telescope has been operating nominally. A new guide camera for the echelle was installed, and only minor issues with it are remaining. DIS had a monthly servicing. The blue camera contamination reappears very shortly after vacuum servicing, whereas the red camera seems to be holding steady and not much has been done to it. In recent discussions the engineering staff concluded that we are reaching a point of diminishing returns on the regular servicing of DIS Blue because the scattered light gets bad again after ~ 5 days. Although it is too late to indicate this in Q3 proposals since they were already submitted, UC reps were asked to tell their users that if they absolutely require DIS Blue for their proposed Q3 observations they should contact Russet ASAP and she will do her best to schedule them immediately following a servicing.

ARCSAT is currently having a balance issue; it will be addressed when the staff can get to it after handling 3.5m issues. People are still getting data, but they experience image jumps and elongated images. It's probably OK for people doing standard aperture photometry. This is a difficult telescope to balance and we recently added a big heavy finder scope and moved some equipment around. Russet reported that this issue may have a seasonal component to it as well.


Q2 3.5m scheduling:

Because the KOSMOS commissioning was moved from late June into September, some dark time opened up in late June (June 26-28). June 19 is also open (100% moon). Users should request any of this time if they can use it. Russet may not be able to update the June schedule until after she finishes the Q3 schedule.


Q2 ARCSAT scheduling:

We still have a couple of open weeks available and we encourage requests for that time.


Q3 3.5m scheduling:

Everyone sent their requests in on time - thank you! Russet is currently building the Q3 schedule. There may be a bit of a crunch in late September but it's a bit early to tell.


NA2 Baffle Project:

Nancy provided an update on this project on behalf of John Wilson. The contract has been let for the stray light analysis of the new baffle design to be done by Photon Engineering. John estimated that the stray light analysis would be done by late July. He would then modify the baffle's mechanical design based on their recommendations and have a final review in late August prior to fabrication. The fabrication should take 2-3 months, thus we expect delivery/commissioning at APO to be in the November-December timeframe.


KOSMOS Update:

Sarah reported that there is a lot of activity underway in their lab. The big instrument support cart is done and just needs painting. They were experimenting with the etching process for manufacturing the slits, and they settled on slit lengths that are comparable to that of DIS. The next steps involve swapping the detector into the new dewar. Conor has the software well underway; it is roughly halfway completed and he needs to do some testing. They are working on a new electrical control box to use with the new slit viewing camera. KOSMOS is currently scheduled to be received at APO at the beginning of the second week in September and will be commissioned on sky shortly thereafter.


Response to Call for White Papers for 2020 Decadal Survey

Nancy reported that the deadline for Activities, Projects, or State of the Profession Considerations (APC) white papers has been set: it is July 10. She re-contacted several observatory directors whom she contacted 8 months ago to see if they would still like to participate. Nancy will get the NOI and draft document online in a shared document, and she will be asking several people within ARC to help out with it. ARC's position on the role of ~4m class telescopes may also be crystallized after the Strategic Planning meeting, so the plan is to work on this white paper once that meeting is over. UC reps were asked to think about whom from their departments might be good candidates to help with this effort.


Discussion of 25th Anniversary / Strategic Planning meeting in 2019

Nancy reminded the Users Committee of some key information concerning the upcoming Strategic Planning Meeting in June:

  • Dates: June 11-12, 2019
  • Location: Sunspot Astronomy and Visitors Center
  • Lodging: Rooms will be blocked off at The Lodge in Cloudcroft. Nancy (ARC) will be paying for the lodging but attendees still need to book their own rooms, Nancy will receive the bill.
  • Attendance: all ARC members and most leasing partners have identified institutional representatives to attend, which is great!
  • Preparation: all institutions are encouraged to meet and have their internal discussions prior to the meeting. The discussion questions are posted on the meeting web site on the Reference Materials page. UC reps were asked to coordinate with their meeting attendees to ensure that these questions get discussed in their departments prior to the Strategic Planning Meeting.

Nancy will post the finalized agenda by the end of today (Monday June 3). She will also remind people of the spreadsheet where anyone interested in carpooling from/to the El Paso airport can provide their travel information. She will likely have some money leftover after paying for the food and lodging, and she will use the remainder to reimburse people's airfare. Attendees are encouraged to hold off on submitting requests for their institutions to cover their travel expenses until Nancy sees how much money she has leftover and thus how much she can provide to attendees to help defray the cost of their airfare.


ACTION ITEMS:

Open action items from previous meetings:

  • Ben Williams will ask Anjum whether she has any Agile data reduction documentation that she could upload to the wiki. Ben got Anjum's new email address and contacted her about this. Both Ben and Nancy are awaiting responses from other people. If nothing new develops prior to the next call Nancy will remove this from the Action Item list and will discuss alternate ways of getting some Agile data reduction info put onto the wiki. STATUS CLOSED. Neither Nancy nor Ben received any replies so they will work on alternatives for getting this information posted on the wiki. It might be as simple as asking someone to write and post some simple CCD data reduction scripts, although Agile has a filter-specific fringing issue that would need to be addressed. They will look at the schedule of fairly recent Agile users and request their assistance.
  • All UC reps should work with their department members to identify who will attend the Strategic Planning Meeting in June. STATUS: CLOSED. All ARC members and some leasing partners have identified their representatives.
  • Nancy will get an updated timeline on the NA2 baffle project from John Wilson. STATUS: CLOSED. See NA2 Baffle discussion above.
  • Ben will send out a link to KOSMOS info for interested prospective users. STATUS: OPEN. This is intended to be a page where prospective users could learn about the general capabilities of KOSMOS at a glance. Sarah and Kal have been worrying more about getting KOSMOS ready for commissioning, so this remains open. Sarah envisions this to be an entry point with capabilities that will be built up to a full manual that they can link to the APO web site.
  • Nancy will disseminate strategic planning discussion questions for the departments by Friday May 10. STATUS: CLOSED.

New action items from this meeting:

  • All UC reps should notify their users that if their Q3 proposals require DIS Blue for their science, they should contact Russet ASAP and let her know.
  • Nancy will post the final agenda and logistical info for the strategic planning meeting by the end of the day.

Next meeting:

Nancy distributed an availability poll for the UC telecons for the remainder of 2019. Because there is no single day and time that works for everyone, she proposed to alternate between the first Monday and first Tuesday of every month, with all calls beginning at 10:30 am Mountain Time. The proposed schedule is thus:

  • Tue July 2
  • Mon Aug 5
  • Tue Sep 3
  • Mon Oct 7
  • Tue Nov 5
  • Mon Dec 2

If anyone objects to this schedule please let Nancy know by Monday June 10, otherwise she will consider it final and will send out reminders.

All Other Business:

None.


Note: See TracWiki for help on using the wiki.